# Measurements of decay branching fractions of the Higgs boson to hadronic final states at the CEPC

### v2

@@ Abstract
The Circular Electron Positron Collider (CEPC) is a large-scale particle accelerator designed to collide electrons and positrons at high energies. One of the primary goals of the CEPC is to achieve high-precision measurements of the properties of the Higgs boson, facilitated by the large number of Higgs bosons that can be produced with significantly low contamination. The measurements of Higgs boson branching fractions into $b\overline{b}/c\overline{c}/gg$ and $WW^{*}/ZZ^{*}$, where the $W$ or $Z$ bosons decay hadronically, are presented in the context of the CEPC experiment, assuming a scenario with 5600 fb-1 of collision data at a center-of-mass energy of 240 GeV. In this study the Higgs bosons are produced in association with a $Z$ boson, with the $Z$ boson decaying into a pair of muons $(\mu^{+}\mu^{-})$, which have high efficiency and high resolution. In order to separate all decay channels simultaneously with high accuracy, the Particle Flow Network (PFN), a graph-based machine learning model, is considered. The precise classification provided by the PFN is employed in measuring the branching fractions using the migration matrix method, which accurately corrects for detector effects in each decay channel. The statistical uncertainty of the measured branching ratio is estimated to be 0.58% in $H\to b\overline{b}$ final state, and approximately 2%-10% in $H\to c\overline{c}/gg/WW^{*}/ZZ^{*}$ final states. In addition, the main sources of systematic uncertainties to the measurement of the branching fractions are discussed.
@@ 1 Introduction
According to theoretical predictions, the branching fractions for the decay of a 125 GeV Higgs boson into $b\overline{b}$, $c\overline{c}$, $gg$, $WW^{*}$, $ZZ^{*}$ are 57.7%, 2.91%, 8.57%, 21.5% and 2.64%, respectively [8, 9, 10]. The Higgs boson decay into $b\overline{b}$, $WW^{*}$, $ZZ^{*}$ were studied by the ATLAS Collaboration using a 13 TeV $pp$ Run 2 dataset collected at a center-of-mass energy of 13 TeV with a luminosity of 139 fb-1 at the LHC. The branching fractions were measured to be $0.53\pm 0.08$, $0.257^{+0.026}_{-0.024}$, $0.028\pm 0.003$, respectively [11].
The work presented here focuses on the determination of the branching fractions of the Higgs boson decaying into a pair of $b$-quark or $c$-quark, gluons, $WW^{*}$ or $ZZ^{*}$ in associated $Z(\mu^{+}\mu^{-})H$ production, where the $W$ or $Z$ bosons decay hadronically, at the CEPC with a center-of-mass energy of 240 GeV and integrated luminosity of 5600 fb-1. The branching fraction measurements for $H\to b\overline{b}/c\overline{c}/gg/WW^{*}/ZZ^{*}$ will be conducted simultaneously considering the major background sources. Since the dominant decay modes of $WW^{*}$ and $ZZ^{*}$ are hadronic, all the five processes result primarily in final states with jets, making it challenging to distinguish between them. Such difficulty is addressed by employing the Particle Flow Network (PFN) [12], which is used for jet tagging, due to its ability to separate these processes. In contrast with traditional jet tagging methods based on QCD theory, which measure branching fractions channel by channel, PFN achieves separation of all channels in a single implementation with high accuracy.
@@ 2 CEPC detector and simulation samples
The signal and background events are both generated using the Monte Carlo (MC) generator Whizard 1.95 [13] and Pythia6 [14] for the fragmentation and hadronization. The response of the CEPC detector is simulated using a Delphes-based software suite for fast detector simulation [15]. The generated MC datasets are normalized to the expected yields in data with an integrated luminosity of 5600 fb-1 by applying a weight to the events of each process.
In this analysis, Higgs production via $ZH$ process is considered to be the dominant process with $Z$ decaying to a pair of muons and Higgs boson decaying in pairs of $b\overline{b}/c\overline{c}/gg/WW^{*}/ZZ^{*}$ is the signal process. The backgrounds originate from processes with two-fermion and four-fermion final states. The two-fermion background processes include $l\bar{l}$, $\nu\bar{\nu}$ and $q\bar{q}$, referring to final states with leptons (l), neutrinos ($\nu$) and quarks (q). The four-fermion background includes $(ZZ)_{h}$, $(ZZ)_{l}$, $(ZZ)_{sl}$, $(WW)_{h}$, $(WW)_{l}$ and $(WW)_{sl}$, referring to final states with leptons (l), hadrons (h) and semi-leptons (sl). Table 1 presents the cross sections of the signal processes. Table 2 provides a summary of the detailed decay modes of the two-fermion and four-fermion backgrounds along with their cross sections. Single $W$ and single $Z$ four-fermion background [16] are neglected in this analysis.
Table 1: Cross sections for the Higgs production via $ZH$ process, where $Z$ boson decays to a muon pair and the Higgs boson decays to $b\overline{b}/c\overline{c}/gg$ and $WW^{*}/ZZ^{*}$, where the $W$ or $Z$ bosons decay hadronically.
@@
Table 2: Detailed decay modes for two-fermion ($l\bar{l}$, $\nu\bar{\nu}$ and $q\bar{q}$) and four-fermion ($(ZZ)_{h}$, $(ZZ)_{l}$, $(ZZ)_{sl}$, $(WW)_{h}$, $(WW)_{l}$, and $(WW)_{sl}$) backgrounds and their cross sections.
@@ 3 Event selection
$H\to b\overline{b}$ $H\to c\overline{c}$ $H\to gg$ $H\to WW^{*}$ $H\to ZZ^{*}$
Muon pair 94.45% 94.24% 94.17% 94.91% 94.43%
Isolation 91.52% 92.81% 93.37% 93.83% 94.04%
$Z$-mass window 95.34% 95.46% 95.49% 91.89% 94.29%
$H$-mass window 99.73% 99.74% 99.74% 99.05% 99.46%
$|\cos\theta_{\mu^{+}\mu^{-}}|<0.996$ 99.65% 99.65% 99.67% 99.65% 99.64%
Total efficiency 81.90% 82.98% 83.46% 80.77% 82.98%
@@
Muon pair 11.95% 0 0.05% 0.08% 46.20% 18.91% 0.00% 11.02% 0.16%
Isolation 91.63% 0 0.40% 2.60% 74.16% 66.47% 0 96.46% 3.75%
$Z$-mass window 41.28% 0 0 0 66.29% 70.41% 0 30.86% 16.67%
$H$-mass window 6.42% 0 0 0 14.22% 14.80% 0 57.32% 0.35%
$|\cos\theta_{\mu^{+}\mu^{-}}|<0.996$ 93.10% 0 0 0 99.07% 99.24% 0 98.40% 98.94%
Total efficiency 0.27% 0.00% 0.00% 0.00% 3.20% 1.30% 0.00% 1.85% 0.00%
Table 3 presents the event selection efficiencies for various signal and background processes, detailing the efficiency at each selection step relative to the previous requirement. In addition, the total efficiency is defined as the ratio of the number of events satisfying all selection criteria to the total number of events expected from the process considered (signal or background). For signal processes, a high efficiency of over 80% is observed. In contrast, two-fermion background processes, primarily $l\bar{l}$, exhibit a total efficiency of around 0.3% and other contributions are negligible. Four-fermion backgrounds, such as $(ZZ)_{l}$, $(ZZ)_{sl}$ and $(WW)_{l}$, have total efficiencies of 3.2%, 1.3% and 1.9%, respectively, while $(ZZ)_{h}$, $(WW)_{h}$, $(WW)_{sl}$ are found to be negligible.
@@ 4 Modeling with Particle Flow Networks
Based on the selection criteria discussed in Section 3, the training process involves a nine-classification task. The signal includes five distinct Higgs decay channels, while the background contains one two-fermion background class ($l\bar{l}$) and three four-fermion classes ($(ZZ)_{l}$, $(ZZ)_{sl}$, $(WW)_{l}$). During the training procedure, 300,000 events for each process are provided to the model, with data split into training, validation and test sets in an 8:1:1 ratio. The PFN is an end-to-end neural network designed to directly utilize the information of the particles to perform event classification. The training variables include the energy of the particle, momentum, $\phi$ which is the azimuth angle, $cos\theta$ where $\theta$ is the polar angle, particle identification number (PID), and impact parameters including $D_{0}$ and $Z_{0}$, which represent coordinates in cylindrical coordinate system.
@@ 5 The model performance
Figure 4: ROC curves for signal and background processes used in classification. The solid lines are the ROC curves of each process considered, and the dashed lines are the ROC curves of the micro and macro average. The dashed black line represents random classification. As can be seen, AUC vaule for each class is above 0.93, indicating a strong classification of the model.
The Receiver Operating Characteristic Curve (ROC) is a graphical representation of the discriminant power of a classifier model as the threshold is varied. Figure 4 depicts the True Positive Rate (TPR) versus the False Positive Rate (FPR) at various discrimination thresholds. The goal of the training is to maximize the TPR while minimizing FPR; therefore the Area Under the Curve (AUC) value serves as an important metric for evaluating the performance of the model. The area under the ROC curve ranges from 0 to 1, where a value of 1 indicates perfect classification and a value of 0.5 suggests a random classification, indicating that the classifier lacks discriminatory power. As can be seen in Figure 4, the AUC vaule for each class is above 0.93, indicating a strong classification performance and the model’s ability to effectively distinguish between classes.
Figure 5: The distributions of classifier outputs for nine categories are shown. Each histogram represents the probability distribution for processes identified within each category.
The classifier outputs are obtained from a nine-unit layer using the SoftMax function. Considering the category $H\to b\overline{b}$ as an example, the SoftMax function computes nine scores for each event, representing the probability distribution for each process being classified as $H\to b\overline{b}$. As illustrated in Figure 5 (b), in the region where the score exceeds 0.8, 99% of the events correspond to the $H\to b\overline{b}$ signal process, while only 1% of the events originate from the $(ZZ)_{sl}$ background. It can be due to the $Z\to\mu^{+}\mu^{-},Z\to u\bar{u}/d\bar{d}$ processes in the $(ZZ)_{sl}$ background, which have the similar properties with the signal, making the classification more challenging. In addition, the PFN has similar performance in other categories. Furthermore, the PFN demonstrates similar performance across other categories. In order to understand the nine-dimensional more intuitively, the t-SNE algorithm [23] is applied to reduce the dimension of the dataset.
@@
As a non-linear dimension reduction algorithm, t-SNE constructs a similarity matrix and aims to preserve the relationships between data points in both high-dimensional and low-dimensional spaces. The differences in high dimensions are represented as distances in two or three dimensions. As shown in Figure 6, $l\bar{l}$ and $(ZZ)_{sl}$ processes are relatively well separated, while signal process as $H\to c\overline{c}$, $H\to gg$, $H\to WW^{*}$ overlap significantly. In addition, $H\to ZZ^{*}$ process shows similarity to all other signal processes, indicating room for further optimization in model training.
Figure 7: The migration matrix for the 9 classes is shown. The horizontal axis represents the prediction of the model for each event in the test set, while the vertical axis indicates the true labels. The sum of values in each row equals 1.
In supervised learning, the migration matrix is used to compare the classified model’s predictions and the true values. Based on the nine classification task, there are nine reconstructed categories, which refer to the process with the highest score for a given event. In Figure 7, the diagonal elements of the matrix represent the correctly classified rates, indicating the purity of each category, while the off-diagonal elements show the misclassification rates. The sum of values in each row equals 1. The migration matrix reflects the overall high accuracy of the model.

### v3

@@ Abstract
The Circular Electron Positron Collider (CEPC) is a large-scale particle accelerator designed to collide electrons and positrons at high energies. One of the primary goals of the CEPC is to achieve high-precision measurements of the properties of the Higgs boson, facilitated by the large number of Higgs bosons that can be produced with significantly low contamination. The measurements of Higgs boson branching fractions into $b\overline{b}/c\overline{c}/gg$ and $\tau\overline{\tau}/WW^{*}/ZZ^{*}$, where the $W$ or $Z$ bosons decay hadronically, are presented in the context of the CEPC experiment, assuming a scenario with 5600 fb-1 of collision data at a center-of-mass energy of 240 GeV. In this study the Higgs bosons are produced in association with a $Z$ boson, with the $Z$ boson decaying into a pair of muons $(\mu^{+}\mu^{-})$, which have high efficiency and high resolution. In order to separate all decay channels simultaneously with high accuracy, the Particle Flow Network (PFN), a graph-based machine learning model, is considered. The precise classification provided by the PFN is employed in measuring the branching fractions using the migration matrix method, which accurately corrects for detector effects in each decay channel. The statistical uncertainty of the measured branching ratio is estimated to be 0.55% in $H\to b\overline{b}$ final state, and approximately 1.5%-16% in $H\to c\overline{c}/gg/\tau\overline{\tau}/WW^{*}/ZZ^{*}$ final states. In addition, the main sources of systematic uncertainties to the measurement of the branching fractions are discussed.
@@ 1 Introduction
According to theoretical predictions, the branching fractions for the decay of a 125 GeV Higgs boson into $b\overline{b}$, $c\overline{c}$, $gg$, $\tau\overline{\tau}$, $WW^{*}$, $ZZ^{*}$ are 57.7%, 2.91%, 8.57%, 6.32%, 21.5% and 2.64%, respectively [8, 9, 10]. The Higgs boson decay into $b\overline{b}$, $WW^{*}$, $ZZ^{*}$ were studied by the ATLAS Collaboration using a 13 TeV $pp$ Run 2 dataset collected at a center-of-mass energy of 13 TeV with a luminosity of 139 fb-1 at the LHC. The branching fractions were measured to be $0.53\pm 0.08$, $0.257^{+0.026}_{-0.024}$, $0.028\pm 0.003$, respectively [11].
The work presented here focuses on the determination of the branching fractions of the Higgs boson decaying into a pair of $b$-quark or $c$-quark, gluons, $\tau\overline{\tau}$, $WW^{*}$ or $ZZ^{*}$ in associated $Z(\mu^{+}\mu^{-})H$ production, where the $W$ or $Z$ bosons decay hadronically, at the CEPC with a center-of-mass energy of 240 GeV and integrated luminosity of 5600 fb-1. The branching fraction measurements for $H\to b\overline{b}/c\overline{c}/gg/\tau\overline{\tau}/WW^{*}/ZZ^{*}$ will be conducted simultaneously considering the major background sources. Since the dominant decay modes of $WW^{*}$ and $ZZ^{*}$ are hadronic, all the six processes result primarily in final states with jets, making it challenging to distinguish between them. Such difficulty is addressed by employing the Particle Flow Network (PFN) [12], which is used for jet tagging, due to its ability to separate these processes. In contrast with traditional jet tagging methods based on QCD theory, which measure branching fractions channel by channel, PFN achieves separation of all channels in a single implementation with high accuracy.
@@ 2 CEPC detector and simulation samples
The signal and background events are both generated using the Monte Carlo (MC) generator Whizard 1.95 [13] and Pythia6 [14] for the fragmentation and hadronization. The response of the CEPC detector is simulated using a Delphes-based software suite for fast detector simulation [15], according to the performance of the baseline detector in the CEPC CDR [7]. The resolution of impact parameter in the $r\phi$ plane is given as:
$\sigma_{r\phi}=5\oplus\dfrac{10}{p(\mathrm{GeV})\sin^{3/2}\theta}\mathrm{\mu m}.$ (1)
The resolution of particle transverse momenta is given as:
$\sigma_{\frac{1}{p_{T}}}=2\times 10^{-5}\oplus\dfrac{1\times 10^{-3}}{p\sin^{3/2}\theta}\mathrm{GeV}^{-1}.$ (2)
The energy resolution of photons is:
$\dfrac{\sigma_{E}}{E}=0.01\oplus\dfrac{0.16}{\sqrt{E(\mathrm{GeV})}},$ (3)
and that of neutral hadrons is:
$\dfrac{\sigma_{E}}{E}=0.03\oplus\dfrac{0.50}{\sqrt{E(\mathrm{GeV})}}$ (4)
In this analysis, Higgs production via $ZH$ process is considered to be the dominant process with $Z$ decaying to a pair of muons and Higgs boson decaying in pairs of $b\overline{b}/c\overline{c}/gg/\tau\overline{\tau}/WW^{*}/ZZ^{*}$ is the signal process, while inclusive decays of $H\to WW^{*}$ and $H\to ZZ^{*}$ are considered. The backgrounds originate from processes with two-fermion and four-fermion final states. The two-fermion background processes include $l\bar{l}$, $\nu\bar{\nu}$ and $q\bar{q}$, referring to final states with leptons (l), neutrinos ($\nu$) and quarks (q). The four-fermion background includes $(ZZ)_{h}$, $(ZZ)_{l}$, $(ZZ)_{sl}$, $(WW)_{h}$, $(WW)_{l}$, $(WW)_{sl}$, $(SZ)_{l}$, $(SZ)_{sl}$, $(SW)_{l}$, $(SW)_{sl}$, $(mix)_{h}$ and $(mix)_{l}$, referring to final states with leptons (l), hadrons (h) and semi-leptons (sl). Table 1 presents the cross sections of the signal processes. Table 2 provides a summary of the detailed decay modes of the two-fermion and four-fermion backgrounds along with their cross sections.
Table 1: Cross sections for the Higgs production via $ZH$ process, where $Z$ boson decays to a muon pair and the Higgs boson decays to $b\overline{b}/c\overline{c}/gg$ and $\tau\overline{\tau}/WW^{*}/ZZ^{*}$, where the $W$ or $Z$ bosons decay hadronically.
@@
Table 2: Detailed decay modes for two-fermion ($l\bar{l}$, $\nu\bar{\nu}$ and $q\bar{q}$) and four-fermion ($(ZZ)_{h}$, $(ZZ)_{l}$, $(ZZ)_{sl}$, $(WW)_{h}$, $(WW)_{l}$, $(WW)_{sl}$, $(SZ)_{l}$, $(SZ)_{sl}$, $(SW)_{l}$, $(SW)_{sl}$, $(mix)_{h}$ and $(mix)_{l}$) backgrounds and their cross sections.
@@ 3 Event selection
$H\to b\overline{b}$ $H\to c\overline{c}$ $H\to gg$ $H\to\tau\overline{\tau}$ $H\to WW^{*}$ $H\to ZZ^{*}$
Simulated events $1.00\times 10^{6}$ $1.00\times 10^{6}$ $1.00\times 10^{6}$ $3.72\times 10^{5}$ $1.00\times 10^{6}$ $1.00\times 10^{6}$
Muon pair 94.45% 94.24% 94.17% 94.94% 94.91% 94.43%
Isolation 91.47% 92.76% 93.31% 94.47% 93.77% 93.99%
$Z$-mass window 96.28% 96.41% 96.41% 92.95% 93.03% 95.28%
$H$-mass window 99.64% 99.66% 99.65% 98.98% 98.88% 99.36%
$|\cos\theta_{\mu^{+}\mu^{-}}|<0.996$ 99.66% 99.66% 99.66% 99.64% 99.65% 99.65%
Total efficiency 82.59% 83.70% 84.14% 81.95% 81.58% 83.72%
@@
Simulated events $1.20\times 10^{8}$ $3.03\times 10^{7}$ $3.03\times 10^{7}$ $3.00\times 10^{6}$ $1.00\times 10^{7}$ $2.60\times 10^{7}$ $2.50\times 10^{7}$ $2.00\times 10^{7}$ $3.00\times 10^{7}$
Muon pair 11.95% 0 0.05% 0.08% 46.21% 18.91% 0.00% 11.03% 0.16%
Isolation 91.67% 0 0.40% 2.60% 74.09% 66.49% 0 96.46% 3.68%
$Z$-mass window 41.82% 0 0 0 67.68% 71.45% 0 34.48% 17.75%
$H$-mass window 6.55% 0 0 0 14.52% 15.02% 0 57.50% 36.76%
$|\cos\theta_{\mu^{+}\mu^{-}}|<0.996$ 90.62% 0 0 0 98.83% 99.56% 0 98.85% 99.15%
Total efficiency 0.27% 0.00% 0.00% 0.00% 3.32% 1.34% 0.00% 2.09% 0.00%
$(SZ)_{l}$ $(SZ)_{sl}$ $(SW)_{l}$ $(SW)_{sl}$ $(mix)_{h}$ $(mix)_{l}$
Simulated events $8.18\times 10^{7}$ $3.20\times 10^{6}$ $3.49\times 10^{6}$ $1.05\times 10^{7}$ $1.29\times 10^{7}$ $1.17\times 10^{7}$
Muon pair 9.92% 0.02% 0 0.00% 0.00% 29.38%
Isolation 44.68% 0 0 0 0 60.77%
$Z$-mass window 18.46% 0 0 0 0 13.78%
$H$-mass window 31.71% 0 0 0 0 35.94%
$|\cos\theta_{\mu^{+}\mu^{-}}|<0.996$ 90.02% 0 0 0 0 62.79%
Total efficiency 0.36% 0.00% 0.00% 0.00% 0.00% 0.19%
Table 3 presents the event selection efficiencies for various signal and background processes, detailing the efficiency at each selection step relative to the previous requirement. In addition, the total efficiency is defined as the ratio of the number of events satisfying all selection criteria to the total number of events expected from the process considered (signal or background). For signal processes, a high efficiency of over 80% is observed. In contrast, two-fermion background processes, primarily $l\bar{l}$, exhibit a total efficiency of around 0.3% and other contributions are negligible. Four-fermion backgrounds, such as $(ZZ)_{l}$, $(ZZ)_{sl}$ and $(WW)_{l}$, have total efficiencies of 3.3%, 1.3% and 2.1%, respectively, while $(ZZ)_{h}$, $(WW)_{h}$, $(WW)_{sl}$ are found to be negligible.
@@ 4 Modeling with Particle Flow Networks
Based on the selection criteria discussed in Section 3, the training process involves a twelve-classification task. The signal includes six distinct Higgs decay channels, while the background contains one two-fermion background class ($l\bar{l}$) and five four-fermion classes ($(ZZ)_{l}$, $(ZZ)_{sl}$, $(WW)_{l}$, $(SZ)_{l}$ and $(mix)_{l}$). During the training procedure, 300,000 events for each process are provided to the model whose weights are all equal to 1, with data split into training, validation and test sets in an 8:1:1 ratio. The PFN is an end-to-end neural network designed to directly utilize the information of the particles to perform event classification. The training variables include the energy of the particle, momentum, $\phi$ which is the azimuth angle, $cos\theta$ where $\theta$ is the polar angle, particle identification number (PID), and impact parameters including $D_{0}$ and $Z_{0}$, which represent coordinates in cylindrical coordinate system.
@@ 5 The model performance
Figure 4: ROC curves for signal and background processes used in classification. The solid lines are the ROC curves of each process considered, and the dashed lines are the ROC curves of the micro and macro average. The dashed black line represents random classification. As can be seen, AUC vaule for each class is above 0.94, indicating a strong classification of the model.
The Receiver Operating Characteristic Curve (ROC) is a graphical representation of the discriminant power of a classifier model as the threshold is varied. Figure 4 depicts the True Positive Rate (TPR) versus the False Positive Rate (FPR) at various discrimination thresholds. The goal of the training is to maximize the TPR while minimizing FPR; therefore the Area Under the Curve (AUC) value serves as an important metric for evaluating the performance of the model. The area under the ROC curve ranges from 0 to 1, where a value of 1 indicates perfect classification and a value of 0.5 suggests a random classification, indicating that the classifier lacks discriminatory power. As can be seen in Figure 4, the AUC vaule for each class is above 0.94, indicating a strong classification performance and the model’s ability to effectively distinguish between classes.
Figure 5: The distributions of classifier outputs for twelve categories are shown. Each histogram represents the probability distribution for processes identified within each category.
The classifier outputs are obtained from a nine-unit layer using the SoftMax function. Considering the category $H\to b\overline{b}$ as an example, the SoftMax function computes twelve scores for each event, representing the probability distribution for each process being classified as $H\to b\overline{b}$. As illustrated in Figure 5 (b), in the region where the score exceeds 0.8, 99% of the events correspond to the $H\to b\overline{b}$ signal process, while only 1% of the events originate from the $(ZZ)_{sl}$ background. It can be due to the $Z\to\mu^{+}\mu^{-},Z\to u\bar{u}/d\bar{d}$ processes in the $(ZZ)_{sl}$ background, which have the similar properties with the signal, making the classification more challenging. In addition, the PFN has similar performance in other categories. Furthermore, the PFN demonstrates similar performance across other categories. In order to understand the twelve-dimensional more intuitively, the t-SNE algorithm [23] is applied to reduce the dimension of the dataset.
@@
As a non-linear dimension reduction algorithm, t-SNE constructs a similarity matrix and aims to preserve the relationships between data points in both high-dimensional and low-dimensional spaces. The differences in high dimensions are represented as distances in two or three dimensions. As shown in Figure 6, $(WW)_{l}$ and $(SZ)_{l}$ processes are relatively well separated, while signal process as $H\to c\overline{c}$, $H\to gg$, $H\to WW^{*}$ overlap significantly. In addition, $H\to ZZ^{*}$ process shows similarity to all other signal processes, indicating room for further optimization in model training.
Figure 7: The migration matrix for the 12 classes is shown. The horizontal axis represents the prediction of the model for each event in the test set, while the vertical axis indicates the true labels. The sum of values in each row equals 1.
In supervised learning, the migration matrix is used to compare the classified model’s predictions and the true values. Based on the twelve classification task, there are twelve reconstructed categories, which refer to the process with the highest score for a given event. In Figure 7, the diagonal elements of the matrix represent the correctly classified rates, indicating the purity of each category, while the off-diagonal elements show the misclassification rates. The sum of values in each row equals 1. The decays of $H\to WW^{*}$ and $H\to ZZ^{*}$ are considered inclusively, while the hadronic decays can be well separated from non-hadronic decays by the classifier. The migration matrix reflects the overall high accuracy of the model.